### v2

# HLSFactory: A Framework Empowering High-Level Synthesis Datasets for Machine Learning and Beyond

Abstract
Machine learning (ML) techniques have been applied to high-level synthesis (HLS) flows for quality-of-result (QoR) prediction and design space exploration (DSE). Nevertheless, the scarcity of accessible high-quality HLS datasets and the complexity of building such datasets present great challenges to FPGA and ML researchers. Existing datasets either cover only a subset of previously published benchmarks, provide no way to enumerate optimization design spaces, are limited to a specific vendor, or have no reproducible and extensible software for dataset construction. Many works also lack user-friendly ways to add more designs to existing datasets, limiting wider adoption and sustainability of such datasets.
In response to these challenges, we introduce HLSFactory, a comprehensive framework designed to facilitate the curation and generation of high-quality HLS design datasets. HLSFactory has three main stages: 1) a design space expansion stage to elaborate single HLS designs into large design spaces using various optimization directives across multiple vendor tools, 2) a design synthesis stage to execute HLS and FPGA tool flows concurrently across designs, and 3) a data aggregation stage for extracting standardized data into packaged datasets for ML usage. This tripartite architecture not only ensures broad coverage of data points via design space expansion but also supports interoperability with tools from multiple vendors. Users can contribute to each stage easily by submitting their own HLS designs or synthesis results via provided user APIs. The framework is also flexible, allowing extensions at every step via user APIs with custom frontends, synthesis tools, and scripts.
To demonstrate the framework functionality, we include an initial set of built-in base designs from PolyBench, MachSuite, Rosetta, CHStone, Kastner et al.’s Parallel Programming for FPGAs, and curated kernels from existing open-source HLS designs. We report the statistical analyses and design space visualizations to demonstrate the completed end-to-end compilation flow, and to highlight the effectiveness of our design space expansion beyond the initial base dataset, which greatly contributes to dataset diversity and coverage.
In addition to its evident application in ML, we showcase the versatility and multi-functionality of our framework through six case studies: I) Using design space sampling in stage 1 to expand the design space covered form a small base set of HLS designs; II) Demonstrating the speedup from the fine-grained design parallelism backend; III) Extending HLSFactory to target Intel’s HLS flow across all stages; IV) Adding and running new auxiliary designs using HLSFactory; V) Integration of previously published HLS data in stage 3; VI) Using HLSFactory to perform HLS tool version regression benchmarking to identify differences between two versions of the same HLS vendor tool.
Code available at https://github.com/sharc-lab/HLSFactory.
I Introduction
Machine learning (ML) techniques have been widely applied to different electronic design automation (EDA) flows including high-level synthesis (HLS)  [1, 2, 3, 4, 5, 6, 7, 8, 9, 10] for quality-of-result (QoR) prediction, optimization, and design space exploration (DSE). A key enabler to the success of such ML techniques is high-quality datasets, most of which are developed for individual studies e.g., [1, 9, 8]. Some recent works have contributed open-source datasets that can be used by other reseachers, e.g.,  [11, 12, 13, 14, 15, 16].
Despite the great benefits of these datasets, there are still fundamental limitations that hinder their wider adoption for ML applications and FPGA research. First, these datasets are usually small or homogeneous, containing only a subset of previously published HLS benchmarks [17, 18, 19, 12], and frequently consisting exclusively of designs that work with one HLS tool from a single vendor. For example, Spector [11] contains only 9 Intel HLS designs, HLSDataset [13] contains 34 AMD/Xilinx HLS designs, and Rosetta [12] contains 6 AMD/Xilinx HLS designs. Second, because of these separately developed HLS datasets, the designs and intermediate/final tool outputs, which serve as important ML model features, are often reported organized in non-standard ad hoc ways. Some datasets contain only source code [18, 17, 19], some datasets contain only resource usage and end-to-end throughput [12] but no clock frequency or power numbers, while some contain only post-implementation results [16]. HLSyn [16] is a dataset for HLS designs targeted towards predicting design quality of FPGAs. It consists of a wider range of programs and compiler directives, enabling performance optimization of designs. However, existing datasets require huge manual effort and deep domain-specific knowledge for ML practitioners if they need a complete, unified, and larger dataset, where they must execute all related HLS tools on their own to re-collect and organize the needed information. Third, it is challenging for external users who want to extend the existing datasets by contributing their own designs, primarily caused by ad-hoc data formats and missing details when building these datasets (e.g., tool version, target FPGA device, clock frequency, implementation flow settings). The fundamental limitation is, however, not the lack of another complete and rich HLS dataset, but rather the lack of a flexible and extensible framework to enable continuous contributions to a standardized and sustainable dataset.
Therefore, in this work, we introduce HLSFactory, the first framework that takes a principled approach to HLS dataset generation, collection, expansion, and integration, aiming to facilitate a continuous and community-wide effort to contribute to the richest HLS dataset, which will keep expanding easily. HLSFactory boasts the following features:
  • Complete and easily extensible with user inputs at multiple stages. HLSFactory has an end-to-end compilation flow including three main stages: design space expansion stage to elaborate single HLS designs at the source-code level into large design spaces; design synthesis stage to execute HLS and FPGA tools; and data aggregation stage for extracting standardized data organization. HLSFactory uses a modular design that allows users to plug in their own designs and tool flows to the dataset with minimal effort at arbitrary stages.
  • Diverse and comprehensive. The initially included dataset covers a wide variety of HLS designs, containing both simple designs synthesized with AMD/Xilinx and Intel tool flows and complex designs using Xilinx-specific features. In addition, HLSFactory has a novel design space expansion and sampling approach, allowing the generation of many design points from a single HLS design, improving overall design space coverage. Further, HLSFactory has comprehensive data metrics from synthesis to implementation, e.g., HLS synthesis reported resource and latency, and post-implementation resource, timing, power, etc.
  • Reproducible and user-friendly. HLSFactory features push-button ease-of-use to run the entire end-to-end dataset generation workflow, allowing anyone to replicate our generated results, and to easily contribute to the framework and the dataset. Specifically, our framework makes it extremely easy for researchers in the FPGA community to contribute data for various FPGA devices.
  • Multi-purpose. Beyond simply being used to build datasets for ML training, HLSFactory is useful for any task where a large, diverse set of HLS runs is needed. This can include regression testing between various HLS tool vendors and versions or benchmarking published academic HLS tools and techniques. Users are able to explore different research domains quickly due to the aforementioned extensibility provided by our work allowing them to carry out experiments with custom flows and tools.
  • High performance and open-source. HLSFactory is capable of running tool flows, such as HLS synthesis, with a user-tunable, fine-grained multiprocessing backend that allows for automatic parallelism across multiple source datasets of HLS designs. We release HLSFactory as an open-source project available on GitHub. HLSFactory provides an end-to-end framework and an initial set of built-in HLS design sources mentioned in this work. It is released as a Python package (with design sources built into the package installation) along with documentation and demos to help users get started with minimal hassle.
In Sec. II, we first provide background on the prior works in existing HLS benchmarks and datasets. Sec. III introduces our HLSFactory framework detailing the three stages. Sec. IV dives into the implementation of HLSFactory, including how it is configured and extended, and our fine-grained parallelism technique to speed up dataset generation. We then perform several case studies in Sec. V that demonstrate the multi-purpose of the proposed framework.
TABLE I: A comparison of HLSFactory with the existing work. : feature supported;
: feature unsupported; : feature partially supported.
Contributions DB4HLS HLSyn HLSDataset HLSFactory
Benchmark — Polybench
Benchmark — MachSuite
Benchmark — Rosetta
Benchmark — CHStone
Collection — PP4FPGA
Collection — Accelerators (§V-D)
Post-HLS Latency
Post-HLS Resources
Post-HLS Artifacts
Post-Impl. Data
HLS Optimization DSL
Fine-Grained Parallel Builds
Xilinx HLS Support
Intel HLS Support
User Extendable to Other Tools
Programmable API
Open Source
II Related Work
HLS community has multiple standard benchmarks for assessing HLS tools including PolyBench [17], CHStone [18], and MachSuite [19], which in total provide around 67 benchmark designs and are far from sufficient for ML training. Rosetta [12], Dai [2], MLSBench [15], DB4HLS [14], HLSDataset [13], and Spector [11] are all recently proposed HLS datasets, where the former four use AMD/Xilinx tools and the last uses Intel tools. MLSBench provides a sampling from different combinations of directives (pragmas) on top of CHStone and MachSuite. DB4HLS provides exhaustive design exploration on 39 designs from MachSuite with a domain-specific language (DSL) for DSE and parallelized synthesis runs. HLSDataset aims to cover all four commonly used benchmarks (PolyBench, CHStone, MachSuite, Rosetta) with a DSL for specifying the design space to sample from. They also illustrate two ML-based case studies for post-implementation resource and power prediction. HLSyn [16] uses control data flow graphs (CDFGs) of compiled HLS kernels for QoR prediction using graph neural network approaches; their designs are sampled from PolyBench and MachSuite. The features of selected prior works and HLSFactory are shown in Table I.
While existing HLS datasets serve as a solid foundation for empowering ML in HLS, they are inherently limited. First, each dataset covers only a subset of commonly used HLS benchmarks, employing ad-hoc data organization, synthesis tools, configurations, and reported metrics, lacking standardization. This fragmentation makes it exceedingly difficult for ML practitioners to effectively utilize all available datasets for training without significant efforts in data reorganization and tool re-execution. Consequently, the quality of ML models is compromised, impeding the advancement of ML in HLS. Second, the lack of standardized data organization and metric reporting poses challenges to dataset extensibility and long-term sustainability, hindering broader user contributions to HLS datasets.
Therefore, rather than introducing yet another HLS dataset, the ML for HLS community urgently requires a standard, extensible, and user-friendly framework. Such a framework would streamline the collection, generation, elaboration, synthesis, and organization of HLS designs and data from diverse sources and community users. This would facilitate the long-term maintenance and expansion of HLS datasets. The pressing need for such a solution is the driving force behind our proposed HLSFactory.
III HLSFactory Framework
Figure 1: A complete overview of the HLSFactory framework with three stages and three entry points where users can contribute their own designs.
III-A HLSFactory Overview
As depicted in Fig. 1, HLSFactory is composed of three stages in its end-to-end synthesis and data extraction flow; before each stage, there is an entry point where users can submit designs and data.
Stage ➊ is design space expansion, aiming at expanding a single HLS design into multiple designs by enumerating different combinations of optimization directives (pragmas), which can significantly increase the number of data points for ML. In this stage, users can submit one or more HLS designs with possible DSE configurations and HLSFactory will extrapolate and expand the design space via a frontend. This frontend features multi-vendor support. Next, HLSFactory will apply a design space sampling to reduce the number of designs to be synthesized by HLS and implementation tools (e.g., Vivado), if needed, to shorten the execution time. We will showcase this usage in Section V-C.
Stage ➋ is design synthesis stage, where vendor-specific HLS and implementation tools are invoked to synthesize HLS designs into RTL code and then placed and routed. In this stage, users can submit their HLS designs to be directly synthesized without extrapolating. We will showcase this usage in Section V-D.
Stage ➌ is data aggregation, where statistics and artifacts are collected from the implemented designs and compiled into a tool-agnostic format for use by downstream tasks such as ML training and benchmarking. In this stage, users can submit their synthesized post-implementation results or datasets to be merged. We will showcase this usage in Section V-E.
III-B Stage 1: Design Space Expansion and Sampling
This stage aims at expanding a single HLS design into multiple by enumerating combinations of optimization directives (either inline or in a separate file), such as loop unroll factors, array partitioning schemes, and whether to pipeline loops. Such expansion is critical for ML usage because of two reasons. First, the original HLS designs and benchmarks are far from sufficient for ML training, and obtaining additional HLS designs is challenging. Second, a key application of ML for HLS is to help designers choose the best optimization directives for their HLS designs by predicting post-HLS-synthesis and post-implementation metrics from HLS source code and directives (e.g., [13, 16, 20, 21]). Therefore, an ML-ready HLS dataset must provide wide coverage of how different choices of optimization directives can impact a design. Note that the design space expansion is expected to be across multiple vendors, tools, and devices.
On the other hand, the expanded design space can be huge, and synthesizing and implementing each design may be prohibitively time-consuming. Therefore, design space sampling is needed.
We define the concept of a frontend pass, which lowers an HLS abstract design to a certain number of concrete designs. An HLS abstract design is not directly synthesizable but contains parameterized directives that require preprocessing. An HLS concrete design is a copy of the abstract HLS design and is augmented with one possible combination of optimization directives from the design space.
III-B1 Vendor-agnostic OptDSL Frontend
loop_opt,3,2
0,lp2,pipeline,unroll,[1 2 4 8]
1,lp3,pipeline,unroll,[1 2 4 8]
2,lp3,,unroll,[1 2 4 8]
set_directive_unroll -factor [factor] k2mm/[name]
set_directive_pipeline k2mm/[name]
Figure 2: A snippet demonstrating the OptDSL syntax.
For design space expansion, all possible combinations of optimization directives for a certain HLS design must be explicitly specified. We propose a frontend using a domain-specific language (DSL), named OptDSL, to specify the design space using a DSE configuration file. Fig. 2 shows an example of the OptDSL syntax, which specifies the choices for how to pipeline or unroll two loops lp2 and lp3.
OptDSL is vendor-agnostic but based on a modified version of a Vitis HLS Tcl script, minimizing the learning curve for designers already accustomed to writing scripts for Vitis HLS. The main feature of OptDSL is the bracket notation that parameterizes an optimization directive with multiple choices. The overall design space is the Cartesian product of the choices for each parameterized directive.
III-B2 Vendor-specific Concrete Design Generation
While abstract designs can be vendor-agnostic, concrete designs are vendor-specific. I.e., different vendor tools have different HLS syntax and directive formats; therefore, during the lowering process, the frontend needs vendor-specific logic to target different tool flows, as depicted in Fig. 1 stage 1. HLSFactory currently provides support for AMD/Xilinx and Intel flows, while other vendors can be easily supported.
The OptDSL file is provided within the abstract design as a file named opt_template.tcl. To lower the abstract design for AMD/Xilinx tools, we generate opt.tcl, a version of opt_template.tcl with bracketed parameters replaced with different concrete values for each design point.
Once these bracketed parameters are substituted, the OptDSL script becomes a valid Tcl script that can be used directly with Vitis HLS.
To support other vendors, the frontend can parse the OptDSL file and identify the specific optimization directives used within it together with their parametrizations. If the provided OptDSL is not sufficient to describe a desired DSE, HLSFactory provides the necessary infrastructure to allow users to specify their own entirely custom frontend as Python code, as long as it conforms to the specified API interface (to be discussed in Sec. IV-A). For instance, a new frontend pass can easily be introduced to parameterize constants in the HLS source code itself: simply copy the existing OptDSL frontend and modify the templating logic and syntax to work with files other than opt_template.tcl.
III-B3 Design Space Sampling
The design space created by the parameterized optimization directives may be extremely large for even a single design, growing exponentially with the addition of each directive. Therefore, it is almost impossible to enumerate every possible design point in the specified design space and execute synthesis and implementation.
HLSFactory natively supports random sampling of design points from the Cartesian product of all combinations of optimization directives. Users can specify the number of sampled design points, trading off design space coverage for dataset build time and storage.
In the future, HLSFactory can be extended to support user-customizable heuristics for selecting design points, utilizing expert knowledge to determine which combinations of optimizations are more useful to sample from and which combinations may result in invalid or redundant designs. For example, the sampling stage can be combined with active learning to determine meaningful design points to be synthesized.
III-C Stage 2: Design Synthesis
The second stage of HLSFactory synthesizes and implements each concrete HLS design, a process we collectively refer to as the design synthesis. This stage also has an entry point for user input—vendor-specific concrete designs can be provided directly at this point without going through design space expansion. This is useful for easy integration of third-party HLS designs where parametrization of the design space may be difficult or unnecessary.
Design synthesis is broken down into two steps: (1) HLSSynth, where an HLS design is synthesized to RTL code, and (2) HLSImpl, where the resulting RTL code is implemented, resulting in a fully placed-and-routed design. For AMD/Xilinx designs, Vitis HLS is used for HLSSynth and Vivado for HLSImpl. However, any vendor tool can easily be integrated into the HLSFactory framework, for example, Yosys [22] or Intel HLS (to be demonstrated in Sec. V-C), by providing Python code for the desired ToolFlow subclasses.
III-D Stage 3: Data Extraction and Aggregation
Once all the frontends and tool flows have been executed on a pool of designs, relevant design data must be extracted and aggregated into structured formats.
HLSFactory provides DataAggregator classes to package HLS synthesis data (estimated latency, resource usage), post-implementation data (timing, resource, and power data), tool execution metadata (version, runtime), and build artifacts (LLVM IR, IP blocks) into shareable datasets.
Furthermore, as in stage 2, users may want to provide input directly at this stage, e.g., when integrating pre-generated data from prior works, where the build process is not reproducible and thus an earlier entry point cannot be used. Therefore, HLSFactory provides an entry point to the data aggregation stage. This entry point can accept fully synthesized and implemented designs, from which HLSFactory’s built-in data aggregators can extract the relevant data, or pre-generated metrics in whatever form is available, which can be used with a custom DataAggregator subclass to adapt such metrics into HLSFactory’s standard output format.
IV Implementation and Usage
IV-A Vendor Agnostic User API
datasets: DesignDatasetCollection = {
"polybench_xilinx": dataset_polybench_builder(WORK_DIR),
"machsuite_xilinx": dataset_machsuite_builder(WORK_DIR),
"chstone_xilinx": dataset_chstone_builder(WORK_DIR),
opt_dsl = OptDSLFrontend(WORK_DIR, random_sample=True,
random_sample_num=N_RANDOM_SAMPLES)
hls_synth = VitisHLSSynthFlow()
hls_impl = VitisHLSImplFlow()
hls_impl_report = VitisHLSImplReportFlow()
datasets_post_frontend = opt_dsl.execute_datasets_parallel(
datasets, n_jobs=N_JOBS)
datasets_post_synth = hls_synth.execute_datasets_parallel(
datasets_post_frontend, n_jobs=N_JOBS)
datasets_post_hls_impl = hls_impl.execute_datasets_parallel(
datasets_post_synth, n_jobs=N_JOBS)
hls_impl_report.execute_datasets_parallel(
datasets_post_hls_impl, n_jobs=N_JOBS)
Figure 3: Example usage of the HLSFactory framework.
HLSFactory is implemented as a Python library and provides a simple user API that allows the framework configuration to be expressed easily as a short Python script (while still allowing for full Python programming if complex configuration is desired).
TABLE II: The HLSFactory User API.
API Functions Description
class Design Single HLS design
class Dataset Multiple HLS designs
class Flow(ABC) Abstract class for arbitrary design flow
Flow.execute(design) Execute a flow on one design
Flow.execute_datasets_parallel(design) Execute a flow on many designs
class Frontend(Flow) Abstract class for frontend design expansion
class OptDSLFrontend(Frontend) Opt DSL frontend for Xilinx HLS designs
class ToolFlow(Flow) Abstract class for EDA tool
class VitisHLSSynthFlow(ToolFlow) Run Vitis HLS synthesis
class VitisHLSImplFlow(ToolFlow) Run Vivado implementation (via Vitis HLS)
class VitisHLSImplReportFlow(ToolFlow) Run Vivado reporting
An example is shown in Fig. 3. The source HLS designs are located and copied to the desired work directory, and the OptDSLFrontend is invoked to sample 10 random design points from each design. The VitisHLSSynthFlow and VitisHLSImplFlow are then be invoked to synthesize and implement each design point, followed by data aggregation using the VitisHLSImplReportFlow to gather data from each implemented design in a standardized format. A full list of the available APIs is available in Table II.
The API also includes abstract base classes (ABCs) that users can subclass to implement their own frontends and tool flows for HLSFactory, for instance, to support another vendor’s HLS tools. HLSFactory abstracts away the complexities of integrating custom user subclasses into the overall dataset generation process, including the use of fine-grained parallelism (to be discussed in Sec. IV-C).
IV-B Directory Structure
Fig. 4 depicts a simple example of the directory structure accepted as input and produced as output of the HLSFactory workflow. As described throughout Sec. III, we first sample the design space for each source abstract design and then run tool flows and data aggregation on the sampled concrete designs. The figure presents the directory structure for the inputs to this process: an abstract design specified in terms of HLS kernel code, an opt_template.tcl file to be used by the OptDSL frontend (described in Sec. III-B1), and auxiliary scripts for the AMD/Xilinx tool flows.
During dataset generation, each abstract design is enumerated into multiple concrete designs, shown in the figure under the newly generated directory source_designs_xilinx__post_frontend. Each concrete design is identified by the concatenation of the name of the original abstract design and a unique hash determined by the combination of optimization directives chosen for that design. This unique combination of optimization directives is generated as the concrete design’s opt.tcl file.
Tool flows and data aggregation run directly within these concrete design directories. After HLS projects are created, synthesized, and implemented (within the hls_prj directory, as depicted), the data aggregation stage collects information from these projects into standardized JSON-formatted files. These JSON files are stored alongside the HLS project directory within each concrete design, making it clear exactly which combination of optimization directives were used to generate the data.
Figure 4: The directory structure that HLSFactory uses. Red are input files; green are the intermediate design points; blue are output files.
IV-C Parallel Build Backend
To build datasets with hundreds and thousands of data points, an efficient backend is needed to dispatch and execute multiple frontend and tool flows in parallel. In the case of HLS, the bottleneck of constructing such datasets is the runtime of the vendor tools themselves. The runtime for synthesizing an HLS design can range from minutes to hours. We may also want to run trial FPGA implementation flows, which can take hours.
To address these needs, every frontend and tool flow component is automatically augmented in a fine-grained parallel build backend based on multiprocessing. Since all frontend and tool flows are based on the abstract base class, we can easily provide this facility to the user. We take advantage of Python’s multiprocessing. We also provide the option to pin each task to its own dedicated CPU core. This approach appears to be a good default to distribute design build workloads on many-core systems.
We also provide a way for users to pool parallelism across dataset collections rather than a single dataset. Users are able to describe a collection of datasets, each with their own set of designs. Instead of dispatching each dataset’s build workloads in its own parallel pool (i.e., naive parallelism), we aggregate all designs into a single parallel pool (i.e., fine-grained parallelism). This feature is automatic for every frontend and tool flow and transparent to the end user.
V Evaluations
We evaluate our work through a series of six case studies which demonstrate HLSFactory’s multifunctionality and ease of use.
V-A Case Study 1: Design Space Coverage
We evaluate how the use of design space expansion in HLSFactory quantitatively and qualitatively improves the overall design space of generated datasets in terms of latency (HLS-reported) and resource usage (post-implementation). In the context of ML, improved design space coverage for these metrics is important for robust model training on downstream tasks, such as ML-based QoR prediction. Thus we perform a case study comparing metrics of the base designs in Polybench, MachSuite, and CHStone ($n=29$) with the designs sampled from them ($n=257$).
Figure 5: Effect of design sampling to cover more design space. Sampled designs cover a wider range of metrics than base designs with no optimizations. Latency is HLS estimated; resources are post-implementation. Note that these are stacked density plots to show the effect of cumulative design sampling.
We start with a quantitative evaluation. Fig. 5 illustrates the cumulative distributions of these metrics as a stacked histogram representing only base designs ($n=29$), half the sampled designs ($n=129$), and all the sampled designs ($n=257$). We highlight that the sampled designs cover a wider range of average-case latency, LUT usage, and FF usage, with denser coverage as $n$ increases. In the case of DSP and BRAM usage, most base designs use none of these resources while sampled designs do.
We then illustrate the qualitative coverage of the design space in Fig. 6. This space is the 2-D embedding space of HLS-reported and post-implementation metrics generated using a PacMAP [23] dimensional reduction. [^1] PaCMAP is one of many dimensional reduction techniques which can be used to visualize data in this manner. However, many dimensional reduction techniques have trade-offs and free parameters with no meaningful defaults. Each of the base designs is depicted as large emphasized points within this embedding space; sampled designs from the same base design (top panel) or the same benchmark (bottom panel) have matching colors. The convex hulls around same-colored points show the portion of the embedding space covered by design space expansion from each base design or benchmark. This clearly shows that sampling from the expanded design space results in non-overlapping coverage that otherwise would not appear in the final dataset.
Figure 6: Embedding of sampled designs across selected benchmarks. Base designs without optimizations are emphasized. Design points and locations are the same between both panels; they are only colored and grouped differently.
V-B Case Study 2: Speedup of Fine-Grained Design Parallelism
We evaluate our fine-grained parallelism strategy described in Sec. IV-C using a case study synthesizing (Vitis HLS) designs sampled from Polybench, MachSuite, and CHStone across 32 cores.
Results are shown in Fig. 7, showing that fine-grained parallelism achieves more than 20% speed up compared with the naive parallelism approach. Such fine-grained parallelism is especially beneficial given the user-specified timeout threshold (annotated as gray bars).
Figure 7: Parallel execution of Vitis HLS synthesis. Top panel shows core utilization over time with naive parallelism across datasets; bottom panel shows our fine-grained design parallelism across datasets.
V-C Case Study 3: Targeting Different Vendors
To demonstrate the extensibility of the first stage of HLSFactory, we show how to add support for Intel’s i++ HLS flow.
As described in Sec. III-B2, HLSFactory includes an OptDSL parser that recognizes Vitis HLS optimization directives in opt_template.tcl, such as the set_directive_unroll and set_directive_array_partition commands. We can therefore build our Intel-lowering frontend on top of this functionality.
Because i++ does not support specifying optimization directives in a separate file, our frontend instead transforms the HLS source code directly to add i++-compatible versions of each directive parsed from the opt_template.tcl file.
While our frontend can often generate exact equivalents for the specified directives, in some cases, i++ has no exact equivalent for a particular directive used by Vitis HLS, such as array_partition directives. In these cases, we substitute similar directives—in this case, a combination of Intel directives hls_numbanks and hls_bankwidth that achieve a similar memory partitioning result.
Since HLSFactory is agnostic to the specific directives being used and does not correlate specific AMD/Xilinx concrete designs with specific Intel concrete designs, directives need not match one-to-one. There is no impact on correctness; substituting similar directives still improves the diversity of the dataset.
Figure 8: Distribution of post-implementation metrics for PolyBench and MachSuite designs ($n=1340$) using Intel’s HLS flow.
In total, our end-to-end Intel flow extends the HLSFactory user APIs in Table II with three Intel equivalents: OptDSLFrontendIntel(Frontend) as described, IntelHLSSynthFlow(ToolFlow) to invoke i++ for HLS, and IntelQuartusImplFlow(ToolFlow) to invoke Quartus for implementation. We run this flow on designs sampled from PolyBench and MachSuite and plot the resulting metrics in Fig. 8. Intel’s HLS tool does not report overall latency estimates, but it optimizes each kernel’s throughput by maximizing clock speed, which we use as a proxy for performance.
V-D Case Study 4: Adding Auxiliary Design Collections
Third-party researchers may have existing, synthesizable, vendor-specific HLS designs to integrate into HLSFactory, but they may not want or need to create an OptDSL specification for them.
For instance, the authors of LightningSim [24] collect 33 synthesizable open-source designs for AMD/Xilinx Vitis HLS to evaluate their simulation tool, including designs from AMD/Xilinx sample code repositories [25, 26], algorithm implementations from Kastner et al.’s Parallel Programming for FPGAs [27], and graph neural network implementations from FlowGNN [28]. These designs are all provided in a standard format, each having a Tcl script setup.tcl to set up a Vitis HLS project for synthesis.
Using the entry point at the design synthesis stage, one graduate student was able to integrate all of these designs into HLSFactory in less than one hour. To match the input directory structure in Fig. 4, we only needed to copy setup.tcl to dataset_hls.tcl with csynth_design appended (HLSFactory’s VitisHLSSynthFlow expects it to setup the project and run synthesis) and add a four-line script dataset_hls_ip_export.tcl to invoke implementation from Vitis HLS. Since we used the entry point after design space expansion, these were concrete designs, not abstract designs, so no opt_template.tcl was required.
Many other works [29, 30, 31, 32] were also easily integrated with HLSFactory in a similar fashion; the code is available online.
V-E Case Study 5: Integrating Released Data from Other Work
Figure 9: Distribution of HLS-estimated metrics from selected HLSFactory benchmarks (PolyBench, CHStone, and MachSuite; $n=167$) vs. HLSyn ($n=3371$).
We may still want to incorporate previously published data have published to build a more comprehensive HLS dataset. HLSFactory’s data aggregation step provides an entry point to incorporate external data sources into our dataset with ease.
We illustrate how HLSFactory can integrate pre-generated data from prior works—in this case, HLSyn [16]. HLSyn provides both the source code (with places to template optimization directives) of their selected kernels, as well as associated metrics for HLS-reported resource usage and latency for sampled designs. We write a DataAggregator subclass to integrate this data into HLSFactory.
The results are illustrated in Fig. 9, showing the distributions of reported HLS metrics sourced from the listed valid designs of HLSyn and a small sampled subset of designs from our base PolyBench, CHStone, and Machsuite datasets.
The HLSyn flow is built on top of AutoDSE [33, 34] and the Merlin compiler [35, 36], both of which are open-source software tools aimed at extending high-level synthesis capabilities for optimized design space exploration (DSE) and source-to-source translations of HLS designs. These tools suggests future work to integrate AutoDSE and the Merlin compiler as custom flows in HLSFactory, allowing designs to be built from the design space specifications defined in AutoDSE and synthesized with the Merlin compiler.
V-F Case Study 6: Regression Benchmarking of HLS Tools
New versions of HLS vendor tools are periodically released and improve both the tool performance (e.g., faster synthesis) and the QoR of synthesized designs (e.g., less resource usage). However, quantifying such improvements across different tool versions is difficult without a way to benchmark a wide range of designs, similar to the regression testing used in traditional software development.
We demonstrate that HLSFactory streamlines regression testing on HLS tools. We compare Vitis HLS versions 2021.1 and 2023.1 using designs sampled from Polybench, Machsuite, and CHStone (with 16 samples per base design). We collect paired samples by synthesizing the same design with both tool versions.
This experiment was set up in a fully self-contained Python script and HLSFactory enabled this initial study to be completed by one graduate student in three hours.
The results are shown in Fig. 10. We show distributions of the tool runtime, HLS-estimated latency, LUT usage, and FF usage across tool versions. We also report the $p$-value for a paired two-tailed Wilcoxon signed-rank test [37, p. 350] [^2] We use the paired Wilcoxon signed-rank test because we cannot assume our metrics are normally distributed. The use of this test is often misrepresented in statistical education with discussion outside the scope of this work [38, 39, 40]. This test compares the “pseudo-medians” of two distributions; we also do not adjust for differences in their variances. We include statistical testing mainly to show how HLSFactory can be used for more in-depth statistical analysis alongside hardware design benchmarking. and indicate cases with a $p$-value less than $\alpha=0.05$ with an asterisk, indicating a statistically significant difference. Note that for certain metrics, the mean and median shift in opposite directions between tool versions.
Figure 10: Distribution of HLS tool metrics from two versions of Vitis HLS.
VI Conclusion
HLSFactory brings a much-needed principled approach to generating datasets of HLS designs. Our case studies show a small sample of what can be done when we have a flexible, reproducible way to generate data from HLS designs; we hope that our brief explorations show that there is substantial untapped potential for future research into how ML can be applied to HLS.
We also consider directions for future extensions of HLSFactory. Our framework currently has no support for collecting post-simulation metrics like vector-based power analysis or simulated latency. Introducing simulation to HLSFactory, particularly for designs where only a high-level C testbench is available rather than an RTL testbench, is a valuable direction for future work.
We hope that, through open-source, this work invites the research community to collaborate and contribute more designs and tool flows and accelerate ML research for EDA applications.